Question 2
Juanita has an opportunity to invest in her friend's clothing store. The initial investment is $10,200 and the expected annual cashflows thereafter are {$300; $700; $1,300; $2,000; $2,000; $5,000; $5,000}. What is Juanita's IRR on this investment? (Allow two decimals in the percentage but do not enter the % sign.)
1
Expert's answer
2020-11-17T07:22:45-0500

The given information can be entered in the table as shown below.




IRR lies between 9"\%" and 10"\%" . By simple interpolation IRR = "9""\%""+""1""\%""\times""\frac{101}{376+101} = 9.21"